OpenStudy (anonymous):

Help for medal!! Select the equation of the line in slope-intercept form: 4x + 9y = 36

OpenStudy (anonymous):

OPTION 1: \[y=-\frac{ 4 }{ 9 }x+4\] OPTION 2: \[y=\frac{ 9 }{ 4 }x+4\] OPTION 3: \[y=-\frac{ 4 }{ 9}x-4\] OPTION 4: \[y=-\frac{ 9 }{ 4 }x-4\]

OpenStudy (anonymous):

@princeharryyy

OpenStudy (princeharryyy):

9y = -4x + 36 divide it by 9 y = -4/9 x + 4 (OPTION 1)

OpenStudy (anonymous):

Thank you @princeharryyy :)

OpenStudy (princeharryyy):

:)

OpenStudy (anonymous):

can you help me with another one @princeharryyy? is the slope of this linear equation, 3x + y = 5, OPTION 1? OPTION 1: 5 OPTION 2: 3 OPTION 3: -3 OPTION 4: -5

OpenStudy (princeharryyy):

OPTION 3: -3

OpenStudy (princeharryyy):

compare it with y = mx + c where m is the slope and c is some constant